OpenStudy (anonymous):

C = 5/9 (F-32) I need to solve for F

OpenStudy (anonymous):

looks like celsius to fahrenheit lol (9/5)C = F - 32 so F = (9/5)C + 32

OpenStudy (anonymous):

ok I see what you did thank you